# Sharpness of Minima in Deep Matrix Factorization: Exact Expressions

### v2

@@ 4 Overparameterized Deep Matrix Factorization
We also provide a self-contained proof of this corollary in Appendix B.2.
Figure 1: Contour map of the loss landscape around a minimum in a 15-layer overparameterized scalar factorization of a random scalar. GD with different step sizes $\eta$, indicated by different colors, is initialized within a radius of $10^{-9}$ from the minimum, in the direction of the Hessian eigenvector corresponding to the largest eigenvalue. The vector $\bm{v}_{1}$ denotes the eigenvector of the Hessian corresponding to the largest eigenvalue, while $\bm{v}_{N}$ denotes the eigenvector corresponding to the smallest eigenvalue (see Appendix C.1). The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Theorem 20.
@@
Figure 2: GD dynamics with different step sizes, $\eta\geq 2/\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$, indicated by different colors, are initialized within a radius of $10^{-15}$ from the minimum in the direction of the Hessian eigenvector corresponding to the largest eigenvalue, for depth-22 matrix factorization, $\bm{M}=\bm{L}\bm{R}^{\top}$, of a random Gaussian matrix, where $\bm{L}\in\mathbb{R}^{10\times 20}$ and $\bm{R}\in\mathbb{R}^{20\times 20}$. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Corollary 40.
@@ 5 Experiments
For the experiment, we first generate the layer dimensions randomly and then construct the optimal layers $\bm{W}_{1}^{*},\bm{W}_{2}^{*},\dots,\bm{W}_{L}^{*}]$ as Gaussian random matrices, with each entry sampled from $N(0,1)$ according to the generated dimensions. Then, we compute $\bm{M}$ or $m$ by $\prod_{j=1}^{L}\bm{W}^{*}_{j}$.
The linear stability analysis relies on a quadratic approximation of the loss function in the vicinity of a global minimum (Wu et al., 2018). To observe whether the escape phenomenon occurs, GD must be initialized at a locally attractive point. However, making a well-informed guess of such an initial point would require computing the third-order characteristics of the loss function, which is not feasible, since the third-order terms describe how rapidly the Hessian changes. Therefore, we initialize GD extremely close to the minimum—on the order of $10^{-15}$ to $10^{-9}$—to ensure that it is initialized at a locally attractive point.
@@
To measure the distance between the convergence point and the minimizer, we plot the normalized $\ell^{2}$-norm of $\bm{w}_{k}-\bm{w}^{*}$ at each iteration. Furthermore, as shown in Fig. 1(a), Figs. 2(a), 3(a) and Figs. 2(b), 3(b), if $\eta>2/\lambda_{\max}$, where $\lambda_{\max}:=\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$, GD always escapes from the minimum, regardless of how close it is initialized to that point. On the other hand, if $\eta=2/\lambda_{\max}$ then GD converges as shown in Figs. 1-3. A catapult in the training error indicates GD’s escape from the basin of a minimum, after which it eventually converges to another minimum as observed by Wu et al. (2018). For additional experiments, see Appendix C.2.
We choose the perturbation direction for the initial point to be the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue, so as to avoid the manifold formed by minimizers as observed in Fig. 2(c). For the methodology used to generate contour maps of the loss landscape near the minimum and to track the trajectories of GD, see Appendix C.1.
6 Conclusion and Discussion
In this paper, we derived an exact expression for the maximum eigenvalue of the Hessian of the squared-error loss for deep matrix factorization problems for any global minimizer. Our experiments demonstrated that the escape phenomenon explored by Wang et al. (2022) in a one-dimensional problem also occurs in deep matrix factorization. Our results also directly extend to deep linear neural network training problems. Furthermore, the results of this paper provide a step towards understanding the implicit biases of gradient-based optimization methods for non-convex problems. It has been empirically observed that explicit regularization prevents the edge-of-stability phenomenon (Liang et al., 2025, Figure 3). Thus, it would be interesting to derive an exact expression for the maximum eigenvalue of the Hessian for the $\ell^{2}$-regularized deep matrix factorization problem.
@@ A.1 Proof of Lemma 10
By using the substitution of variables as if in (5),
@@ A.2 Proof of Lemma 13
Then, the directional second derivative $\nabla^{2}f(\bm{W}_{1},\cdots,\bm{W}_{L})[\bm{U}_{1},\cdots,\bm{U}_{L}]$ equals to
@@
Appendix B Proofs from Section 4
B.1 Proof of Theorem 5
@@
Note that $\sigma_{\max}(\bm{K}^{\top}\bm{K})=\sigma_{\max}(\bm{K}\bm{K}^{\top})$. Note that for any two block matrices $\bm{A}$ and $\bm{B}$ such that
@@ B.2 Proof of Corollary 40
Since $\sigma_{\max}{(\bm{L}^{*})}\norm{\bm{u}}_{2}+\sigma_{\max}(\bm{R}^{*})\norm{\bm{v}}_{2}\geq 0$ $\forall\bm{u},\bm{v}:\norm{\bm{u}}_{2}^{2}+\norm{\bm{v}}_{2}^{2}=1$, we can write the following equivalence.
$\mathop{\mathrm{arg\,max}}_{\begin{subarray}{c}\bm{u},\bm{v}\\ \|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}=1\end{subarray}}2\bigl(\sigma_{\max}(\bm{L}^{*})\|\bm{u}\|_{2}+\sigma_{\max}(\bm{R}^{*})\|\bm{v}\|_{2}\bigr)^{2}=\mathop{\mathrm{arg\,max}}_{\begin{subarray}{c}\bm{u},\bm{v}\\ \|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}=1\end{subarray}}\sigma_{\max}(\bm{L}^{*})\|\bm{u}\|_{2}+\sigma_{\max}(\bm{R}^{*})\|\bm{v}\|_{2}.$ (85)
This constrained optimization problem is a specific case of the following general constrained optimization problem.
$\min_{\bm{x}\in\mathbb{R}_{+}^{2}}-\bm{c}^{\top}\bm{x}\quad\text{s.t}\quad\norm{\bm{x}}_{2}^{2}\leq 1,$ (86)
where $\bm{c}\geq\bm{0}$. Notice that this is a convex optimization problem. Moreover, it is straightforward to verify that the optimal solution must lie on the boundary of the constraint. To solve it, we can formulate the Lagrangian as follows:
$L(\bm{x},\mu)=-\bm{c}^{\top}\bm{x}+\mu(\bm{x}^{\top}\bm{x}-1)\quad\mu\geq 0.$ (87)
We know that optimal solution satisfies KKT conditions. Therefore,
$\bm{c}+2\mu\bm{x}^{*}=0\rightarrow\bm{x}^{*}=\frac{\bm{c}}{2\mu}\rightarrow\bm{x}^{*}=\frac{\bm{c}}{\norm{\bm{c}}_{2}}.$ (88)
If you select $\bm{c}=\begin{bmatrix}\sigma_{\max}(\bm{L}^{*})\\ \sigma_{\max}(\bm{R}^{*})\end{bmatrix}$ then $\bm{x}^{*}=\begin{bmatrix}\frac{\sigma_{\max}(\bm{L}^{*})}{\sqrt{\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}}}\\ \frac{\sigma_{\max}(\bm{R}^{*})}{\sqrt{\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}}}\end{bmatrix}$. This implies
$\max_{\begin{subarray}{c}\bm{u},\bm{v}\\ \|\bm{u}\|_{2}^{2}+\|\bm{v}\|_{2}^{2}=1\end{subarray}}2\Bigl(\sigma_{\max}(\bm{L}^{*})\|\bm{u}\|_{2}+\sigma_{\max}(\bm{R}^{*})\|\bm{v}\|_{2}\Bigr)^{2}=2(\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}).$ (89)
Therefore,
$\lambda_{\max}\!\bigl(\nabla^{2}\mathcal{L}(\bm{L}^{*},\bm{R}^{*})\bigr)=\max_{\begin{subarray}{c}\bm{U},\bm{V}\\ \|\bm{U}\|_{F}^{2}+\|\bm{V}\|_{F}^{2}=1\end{subarray}}2\bigl\|\bm{L}^{*}\bm{U}^{\top}+\bm{V}\bm{R}^{*\top}\bigr\|_{F}^{2}\;\leq\;2\bigl(\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}\bigr).$ (90)
@@
Appendix C Additional Experimental Results
C.1 Visualization of the Contour Map of the Loss Landscape

### v3

@@ 4 Overparameterized Deep Matrix Factorization
We also provide a self-contained proof of this corollary in Appendix C.2. This result was also recently observed by Josz (2025) independently.
5 Flatness
In this section, we reveal remarkable aspects of the loss landscape of general deep matrix factorization problems. First, we show that an optimal solution in deep overparameterized scalar factorization problem is flat if and only if the product of spectral norms of left and right intermediate networks is constant across layers. This is a direct consequence of Theorem 21.
Corollary 7.
Consider the following objective function
$\mathcal{L}(\bm{w}):=({m-\bm{w}_{L}\bm{W}_{L-1}\cdots\bm{W}_{2}\bm{w}_{1}})^{2},$ (29)
where $m\in\mathbb{R}$, $d_{0}=d_{L}=1$, $\bm{w}_{L}\in\mathbb{R}^{1\times d_{L-1}}$ and $\bm{w}_{1}\in\mathbb{R}^{d_{1}\times 1}$. For hidden factors (layers), i.e, for all $i\in\{2,3,\cdots,L-1\}$, we have $\bm{W}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$. Then, $\bm{w}^{*}\in\Omega_{F}$ if and only if
$\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=|m|^{1-\frac{1}{L}}\quad\forall k\in[L],$ (30)
Proof.
For a minimizer $\bm{w}^{*}$, let us assume that $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=|m|^{1-\frac{1}{L}}$ for all $k\in[L]$. Then, by Theorem 21, $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))=2L\times m^{2(1-1/L)}$. This implies that $\bm{w}^{*}$ is flat (Mulayoff & Michaeli, 2020, Theorem 1). Now, assume that $\bm{w}^{*}$ is flat. (Mulayoff & Michaeli, 2020, Theorem 2) showed that for any flat minimum $\bm{w}^{*}$,
$\sigma_{\max}(\bm{A}_{k})=|m|^{1-\frac{k}{L}}\quad\text{and}\quad\sigma_{\max}(\bm{B}_{k})=|m|^{\frac{k-1}{L}}\quad\forall k\in[L].$ (31)
Hence, $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=|m|^{1-\frac{1}{L}}$ for all $k\in[L]$. ∎
Second, we show that being flat in depth-2 matrix factorization is equivalent to being spectral-norm balanced. This result is a direct consequence of Corollary 25 and AM-GM inequality (Josz, 2025, see also Lemma 5).
Corollary 8.
Consider the following depth-$2$ matrix factorization objective
$\mathcal{L}(\bm{L},\bm{R})=\norm{\bm{M}-\bm{L}\bm{R}^{\top}}_{F}^{2},$ (32)
where $\bm{M}\in\mathbb{R}^{d_{L}\times d_{0}}$ is the target matrix and $\bm{L}\in\mathbb{R}^{d_{L}\times k}$, $\bm{R}\in\mathbb{R}^{d_{0}\times k}$ such that $k\geq\min\{d_{0},d_{L}\}$. Then, $(\bm{L}^{*},\bm{R}^{*})\in\Omega_{F}$ if and only if
$\sigma_{\max}(\bm{L}^{*})=\sigma_{\max}(\bm{R}^{*})=\sqrt{\sigma_{\max}(\bm{M})}.$ (33)
Remark 9.
We discuss further implications of this equivalence in the subsequent section.
Now, we show that a minimizer of deep matrix factorization loss is flat if the product of spectral norms of left and right intermediate factors is constant across layers.
Corollary 10.
Consider the optimization objective in (15). If $\bm{w}^{*}\in\Omega$ then
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)\leq 2\sum_{k=1}^{L}\sigma_{\max}(\bm{A}_{k})^{2}\sigma_{\max}(\bm{B}_{k})^{2},$ (34)
where $\bm{A}_{k}=\prod_{i=k+1}^{L}\bm{W}_{i}^{*}$ and $\bm{B}_{k}=\prod_{i=1}^{k-1}\bm{W}_{i}^{*}$ (see Appendix B). This implies that if $\bm{w}^{*}$ satisfies $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=\sigma_{\max}(\bm{M})^{1-\frac{1}{L}}$ for all $k\in[L]$, then $\bm{w}^{*}$ is flat.
Proof.
Mulayoff & Michaeli (2020) showed that for any $\bm{w}^{*}\in\Omega$,
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)\geq 2L\sigma_{\max}(\bm{M})^{2(1-1/L)}.$ (35)
Now, assume that $\sigma_{\max}(\bm{A}_{k})\sigma_{\max}(\bm{B}_{k})=\sigma_{\max}(\bm{M})^{1-\frac{1}{L}}$ for all $k\in[L]$. Then, by (34),
$\lambda_{\max}\bigl(\nabla^{2}\mathcal{L}(\bm{w}^{*})\bigr)=2L\sigma_{\max}(\bm{M})^{2(1-1/L)}.$ (36)
This implies that $\bm{w}^{*}$ is flat (Mulayoff & Michaeli, 2020, Theorem 1). ∎
6 Discussion
In this section, we thoroughly examine the claims made by Ding et al. (2024) and Ghosh et al. (2025) that flat minima coincide with Frobenius-norm balanced minima in depth-2 matrix factorization and then we discuss how misleading sharpness measures can be for the loss landscape analysis.
Strict assumptions. First, we examine the latter work, where Ghosh et al. (2025) examined the learning dynamics of deep linear networks within the deep matrix factorization loss beyond the edge of stability. Their analysis relies on two strict assumptions, which the latter follows the first one. First, they introduce the singular vector stationary set, i.e., for any initialization of GD from this set, GD does not rotate the layers during training. In Proposition 4, they prove that the balanced initialization they considered in the paper, i.e., $\bm{W}_{L}(0)=\bm{W}_{L-1}(0)=\cdots=\bm{W}_{1}(0)=\alpha\bm{I}$ where $\alpha\in\mathbb{R}$, is a member of the singular vector stationary set. This is leveraged to decouple the dynamics of the singular vectors and singular values. Second, for this specific initialization, singular values remain balanced during training, which is not true for arbitrary initialization. Under these assumptions, their optimization objective becomes a deep scalar factorization problem in which the spectral norm equals the Frobenius norm. Therefore, they claim that Frobenius-norm balanced minima correspond to flat minima in deep matrix factorization problem. However, this is only valid from the perspective of GD that initialized at a specific set of points. Globally, we can find minimizers that are flat but not Frobenius-norm balanced. To see this, consider a depth-2 matrix factorization problem, i.e.,
$\min_{\bm{L},\bm{R}}\mathcal{L}(\bm{L},\bm{R})\quad\text{such that}\quad\mathcal{L}(\bm{L},\bm{R})=\norm{\bm{M}-\bm{L}\bm{R}^{\top}}_{F}^{2},$ (37)
where $\bm{M}\in\mathbb{R}^{m\times n},\bm{L}\in\mathbb{R}^{m\times k}$ and $\bm{R}\in\mathbb{R}^{n\times k}$. Suppose $m=n=k=3$ and $\bm{M}=\text{diag}(11,2,1)$. We know that for any flat minimum $(\bm{L}^{\prime},\bm{R}^{\prime})$, $\lambda_{\max}(\nabla^{2}(\mathcal{L}(\bm{L}^{\prime},\bm{R}^{\prime})))=4\times\sigma_{\max}(\bm{M})$, which equals $44$ (Mulayoff & Michaeli, 2020). By definition, any minimum whose sharpness is equal to $44$ is a flat minimum. Let’s investigate a specific minimizer $(\bm{L}^{*},\bm{R}^{*})$ such that $\bm{L}^{*}=\text{diag}(\sqrt{11},\frac{2}{3},1)$ and $\bm{R}^{*}=\text{diag}(\sqrt{11},3,1)$. Using Corollary 25, $\lambda_{\max}(\nabla^{2}(\mathcal{L}(\bm{L}^{*},\bm{R}^{*})))=44$, which means that $(\bm{L}^{*},\bm{R}^{*})$ is a flat minimum. On the other hand, by definition of balancedness, i.e., $\bm{L}\bm{L}^{\top}=\bm{R}\bm{R}^{\top}$, $(\bm{L}^{*},\bm{R}^{*})$ is not a balanced minimizer. Therefore, any flat minimum is not necessarily a balanced minimizer.
Sharpness measures are delusive. Ding et al. (2024) showed that flat minima recover the groundtruth matrix in low rank matrix recovery. Furthermore, they showed that norm-minimal, balanced, and flat solutions coincide in depth-$2$ matrix factorization. They measure sharpness of a minimum by using scaled trace. Under this measure, flat minima coincide with Frobenius-norm balanced minima; however, as we showed above, this is not necessarily true when the worst-case sharpness, i.e., maximum Hessian eigenvalue, is used as the sharpness measure. In fact, we showed that flat minima are spectral-norm balanced when the maximum Hessian eigenvalue is used as the sharpness measure. This means that a loss landscape analysis in general matrix factorization problems using a specific sharpness measure might not lead to the same inferences as those made by using a different metric.
The necessity of an absolute sharpness measure. There is still no consensus in the literature regarding the definition of flatness. For instance, while several works define flat minima as global minimizers that minimize the maximum eigenvalue of the Hessian of the loss (Mulayoff & Michaeli, 2020; Liu et al., 2021; Marion & Chizat, 2024), others define them as global minimizers that minimize the trace of the Hessian (Dinh et al., 2017; Gatmiry et al., 2023). Even a scaled version of the Hessian trace designed to define flat minima in depth-2 matrix factorization problems (Ding et al., 2024), as we mentioned before. As we have shown in Section 5 and discussed above, different sharpness measures can lead to contradictory interpretations of balanced minima. Thus, the necessity of a robust sharpness measure that consolidates the interpretations from existing ones is a matter of urgency. Addressing this concern, Josz (2025) showed how minimizing the maximum Hessian eigenvalue over the solution set can disregard the higher-order variation near minimizers and result in misleading interpretations of flat minima. Therefore, Josz (2025) defined the flat minima of any smooth function $\mathcal{L}$ as the local minima of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}))$ under the constraint $\mathcal{L}(\bm{w})=\mathcal{L}(\bm{w}^{*})$ and demonstrated that this notion coincides with other notions of flatness in the depth-2 matrix factorization.
@@ 6 Discussion
Figure 1: GD dynamics with different step sizes, $\eta\geq 2/\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$, indicated by different colors, are initialized within a radius of $10^{-15}$ from the minimum in the direction of the Hessian eigenvector corresponding to the largest eigenvalue, for depth-22 matrix factorization, $\bm{M}=\bm{L}\bm{R}^{\top}$, of a random Gaussian matrix, where $\bm{L}\in\mathbb{R}^{10\times 20}$ and $\bm{R}\in\mathbb{R}^{20\times 20}$. The vector $\bm{v}_{1}$ denotes the eigenvector of the Hessian corresponding to the largest eigenvalue, while $\bm{v}_{N}$ denotes the eigenvector corresponding to the smallest eigenvalue. The value of $\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$ is computed using the closed-form expression derived in Corollary 25.
@@ 7 Experiments
The dynamical stability analysis relies on how accurate the quadratic approximation of the loss function is in the $\delta$-neighborhood of $\bm{x}^{*}$ (Wu et al., 2018) (also see Definition 1). Therefore, we initialize GD in a $\delta$-neighborhood with $\delta$ on the order of $10^{-15}$ to $10^{-9}$ to ensure validity of the quadratic approximation. To observe escape phenomenon clearly, we choose the perturbation direction for the initial point to be the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue, so as to avoid choosing a direction that is orthogonal to the eigenspace corresponding to the eigenvalues larger than $2/\eta$.
Note that, we could have also considered a random perturbation direction uniformly from $\mathbb{S}^{N-1}$. Mulayoff & Michaeli (2020) showed that for the deep matrix factorization problem (which has the same Hessian structure with deep linear networks), the Hessian is rank-deficient at all minima by at least the order of 1(1/L)1-(1/L). This means that at least 1(1/L)1-(1/L) of the eigenvalues are zero at a minimum. Thus, the probability of choosing a direction that is orthogonal to the eigenspace corresponding to eigenvalues larger than $2/\eta$ is 0. Therefore, random perturbations would lead to the escape phenomenon with probability $1$.
On the other hand, for stable minima, it is important to choose the direction as the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue. The reasoning is the same. If you choose a direction that is not orthogonal to the eigenspace corresponding to the eigenvalues that are zero, then GD never converges to $\bm{w}^{*}$. This means that if we choose the perturbation direction randomly, then with probability 1, we choose a direction that is not orthogonal to the eigenspace corresponding to the eigenvalues that are zero. Therefore, for the experiment, it is convenient to choose perturbation direction as the eigenvector of $\nabla^{2}\mathcal{L}(\bm{w}^{*})$ corresponding to the largest eigenvalue.
@@
To measure the distance between the convergence point and the minimizer, we plot the normalized $\ell^{2}$-norm of $\bm{w}_{k}-\bm{w}^{*}$ at each iteration. Furthermore, as shown in Fig. 1 and Figs. 2(a)-2(c) if $\eta>2/\lambda_{\max}$, where $\lambda_{\max}:=\lambda_{\max}(\nabla^{2}\mathcal{L}(\bm{w}^{*}))$, GD always escapes from the minimum. On the other hand, if $\eta=2/\lambda_{\max}$ then GD converges as shown in Figs. 2(d)-2(f). A catapult in the training error indicates GD’s escape from the basin of a minimum, after which it eventually converges to another minimum as observed by Wu et al. (2018). For the methodology used to generate contour maps of the loss landscape near a minimum, see Appendix D.1, and for additional experiments, see Appendix 3.
8 Conclusion
In this paper, we derived an exact expression for the maximum eigenvalue of the Hessian of the squared-error loss for overparameterized deep matrix factorization problems at any minimizer. We also showed that this expression simplifies considerably for depth-2 matrix factorization and deep, overparameterized scalar factorization. To complement our theory, we conducted GD experiments that crucially rely on our exact expression of the sharpness to observe the escape phenomenon during training. Then, we showed how flat minima are spectral-norm balanced in depth-2 matrix factorization, and we showed that a minimizer of deep overparameterized scalar factorization loss is flat if and only if the product of spectral norms of left and right intermediate factors is constant across layers. Furthermore, we showed that a minimizer of deep matrix factorization loss is flat if the product of spectral norms of left and right intermediate factors is constant across layers. These results led to the following remarkable observation: flat minima are not necessarily Frobenius-norm balanced, which is contradictory to the claim made in (Ghosh et al., 2025). Finally, we discussed how delusive the sharpness measures can be for the loss landscape analysis. Therefore, there is an urgent need for future research to explore a robust sharpness measure that consolidates the interpretations from existing ones.
@@ A.1 Proof of Lemma 11
By using the substitution of variables as in (6),
@@ A.2 Proof of Lemma 14
Then, the directional second derivative $\nabla^{2}f(\bm{W}_{1},\cdots,\bm{W}_{L})[\bm{U}_{1},\cdots,\bm{U}_{L}]$ equals
@@
Appendix B Warm-Up: Deep Overparameterized Scalar Factorization
Proof.
According to (13) and (14),
$=\max_{\begin{subarray}{c}\bm{u}_{1},\bm{u}_{2},\dots,\bm{u}_{L}:\\ \sum_{i=1}^{L}\norm{\bm{u}_{i}}_{2}^{2}=1\end{subarray}}2\Biggl(\sum_{i=1}^{L}\sigma_{\max}\Bigl(\prod_{j=i+1}^{L}\bm{W}^{*}_{j}\Bigr)\sigma_{\max}\Bigl(\prod_{j=1}^{i-1}\bm{W}_{j}^{*}\Bigr)\norm{\bm{u}_{i}}_{2}\Biggr)^{2},$ (65)
where $\bm{U}_{i}\in\mathbb{R}^{d_{i}\times d_{i-1}}$ and $\bm{u}_{i}\in\mathbb{R}^{d_{i}d_{i-1}}$. We can upper bound the right-hand side of (61) by using the triangle inequality. By applying the vectorization trick of the Kronecker product, we can rewrite (62). Then, noting the fact that for any matrix $\bm{A}\in\mathbb{R}^{m\times n}$ and vector $\bm{x}\in\mathbb{R}^{n}$, $\norm{\bm{Ax}}_{2}\leq\sigma_{\max}(\bm{A})\norm{\bm{x}}_{2}$, we can upper bound the right-hand side of (63). Note that for any matrix $\bm{A}$ and $\bm{B}$, $\sigma_{\max}(\bm{A}\otimes\bm{B})=\sigma_{\max}(\bm{A})\sigma_{\max}(\bm{B})$. Hence, we can rewrite (64). Then, by using the Cauchy–Schwarz inequality,
Then, it suffices to show that there exists a direction $\bm{U}_{1}^{*},\bm{U}_{2}^{*},\dots,\bm{U}_{L}^{*}]$ along which the bound in (66) is achieved.
Then,
Since the upper bound is achieved, it implies
Appendix C Proofs from Section 4
C.1 Proof of Theorem 5
@@
Note that $\sigma_{\max}(\bm{K}^{\top}\bm{K})=\sigma_{\max}(\bm{K}\bm{K}^{\top})$, and for any two block matrices $\bm{A}$ and $\bm{B}$ such that
@@ C.2 Proof of Corollary 25
Then, by using Cauchy-Schwarz inequality,
$\lambda_{\max}\!\bigl(\nabla^{2}\mathcal{L}(\bm{L}^{*},\bm{R}^{*})\bigr)\;\leq\;2\bigl(\sigma_{\max}(\bm{L}^{*})^{2}+\sigma_{\max}(\bm{R}^{*})^{2}\bigr).$ (88)
@@
Appendix D Additional Experimental Results
D.1 Visualization of the Contour Map of the Loss Landscape